The equation of motion for a particular mass m suspended from a spring of constant k, is:

x=0.40cos(0.70t-0.30)

where x is the displacement of mass from its rest position, t is the time and the phase angle is in radians

From this equation find:

a)The amplitude

b)The frequency of vibration

c)The period

d)The phase angle in degrees

e)The ratio k/m for the spring


1
Expert's answer
2022-02-11T15:30:37-0500

"x=0.40\cos(0.70t-0.30),"

a)

"A=0.40,"

b)

"\nu=\frac{\omega}{2\pi}=0.11,"

c)

"T=\frac 1\nu=9,"

d)

"\varphi=\frac{180}{\pi}\varphi_0=17°,"

e)

"\frac km=\omega^2=0.49."